# The Energy Cost of Artificial Intelligence Lifecycle in Communication Networks

### v3

@@ VII Energy Cost of Inference
To further quantify this relationship, we first compare the outcomes of Eq. 26, Eq. 28, and Eq. 30, depicted in Fig. 8, where we assume the virtualization is contributing additional $10\%$ to the total energy cost ($\gamma_{\mathrm{v}}=10\%$). The figure shows that the energy consumption per bit of developing a 3 layer - 10 node MLP is $2.41\times 10^{-6}[J/b]$, which is approximately $3.7$ times higher than the energy consumption per bit of one inference ($E_{\mathrm{inf,p,b}}=6.57\times 10^{-7}[J/b]$). Furthermore, we can observe that the energy consumption per bit decreases with more inferences, dropping from $2.11\times 10^{-6}[J/b]$ for $100$ inferences to $7.73\times 10^{-7}[J/b]$ for $1000$ inferences, which is a $2.73$ times of improvement on energy efficiency. Moreover, we can observe that models with architectural blocks that require higher computational complexity need significantly more energy per bit during development compared to simpler models and their energy cost per bit of one inference. For instance, the transformer consumes $3.74\times 10^{-4}\leavevmode\nobreak\ [\mathrm{J/b}]$, which means it is nearly $165$ and $20.7$ times less efficient than the MLP in making one inference. This is also confirmed in Fig. 9. As the number of inferences $\gamma$ in the current model increases, the energy efficiency of the AI model lifecycle improves.
@@ IX Conclusions and Future Work
In this work, we have proposed a novel metric, namely, eCAL. Unlike traditional metrics that focus only on energy required for transmission, computing infrastructure, or AI models separately, eCAL can capture the overall energy cost of generating inferences in communication system during the entire lifecycle of a trained model. We have proposed a detailed methodology to determine the eCAL of an AI enhanced communication system by breaking it down into various data manipulation components, such as data collection, preprocessing, training, evaluation, and inference, and analyzing the complexity and energy consumption of each component. The proposed metric demonstrates that the more a model is utilized, the more energy-efficient each inference becomes. For example, considering a simple MLP architecture, the energy consumption per bit for 100 inferences is $2.73$ times higher than for 1000 inferences. Furthermore, we have developed an open-source, modular energy estimation tool based on our approach. This tool enables fine-grained analysis of energy consumption across components, allowing developers to evaluate and optimize AI energy efficiency in various deployment scenarios.
Through the proposed eCAL metric, this study lays the foundation for understanding the energy consumption of AI-enhanced emerging communication architectures. Future work may focus on extending this study beyond the assumptions, case studies, and configurations provided in this study to cover existing and emerging communication system designs that are increasingly relying on AI techniques for their automation and performance optimization, especially the emerging open radio access network (O-RAN) based cellular systems. The exploration of the relationship between model performance and eCAL for 6G verticals, related network slices, online, distributed, fine tuned and federated model lifecycles may also be worthy of pursuing. Finally, the development of dynamic energy management algorithms in view of a more sustainable operation is highly relevant.
@@ Acknowledgements
This work was supported in part by the HORIZON-MSCA-IF project TimeSmart (No. 101063721), the European Commission NANCY project (No.101096456), and by the Slovenian Research Agency under grants P2-0016 and J2-50071.

### v4

@@ VII Energy Cost of Inference
Figure 13: Energy cost of AI model lifecycle (eCALeCAL) for different models over the number of inferences ($\gamma$), log scale on both axes.
To further quantify this relationship, we first compare the outcomes of Eqs. (27), (29) and (31), depicted in Fig. 12. The figure shows that the energy consumption per bit of developing a 3 layer - 10 node MLP is $2.41\times 10^{-6}[J/b]$, which is approximately $3.7$ times higher than the energy consumption per bit of one inference ($E_{\mathrm{inf,p,b}}=6.57\times 10^{-7}[J/b]$). Furthermore, we can observe that the energy consumption per bit decreases with more inferences, dropping from $2.11\times 10^{-6}[J/b]$ for $100$ inferences to $7.73\times 10^{-7}[J/b]$ for $1000$ inferences, which is a $2.73$ times of improvement on energy efficiency. Moreover, we can observe that models with architectural blocks that require higher computational complexity need significantly more energy per bit during development compared to simpler models and their energy cost per bit of one inference. For instance, the transformer consumes $3.74\times 10^{-4}\penalty 10000\ [\mathrm{J/b}]$, which means it is nearly $165$ and $20.7$ times less energy efficient per bit than the MLP in making $100$ and $1000$ inferences, respectively. Fig. 13 further confirms that as the number of inferences $\gamma$ in the current model increases, the EE of the AI model lifecycle improves.
Implementation guidelines: The open-source eCAL calculator (Python), can be used standalone to compute eCAL values using provided configuration files. Users can study custom scenarios, such as different transmit powers, overheads, or neural network configurations by editing these files. Thanks to its modular design, new protocols or neural architectures can be added via dedicated class implementations. Python-based simulators can integrate it directly by importing its functionality, while other simulators may require minor additional development.
@@ IX Conclusions and Future Work
In this work, we have proposed a novel metric, namely, eCAL. Unlike traditional metrics that focus only on energy required for transmission, computing infrastructure, or AI models separately, eCAL can capture the overall energy cost of generating inferences in communication system during the entire lifecycle of a trained model. We have proposed a detailed methodology to determine the eCAL of an AI enhanced communication system by breaking it down into various data manipulation components, such as data collection, preprocessing, training, evaluation, and inference, and analyzing the complexity and energy consumption of each component. The proposed metric demonstrates that the more a model is utilized, the more energy-efficient each inference becomes. For example, considering a simple MLP architecture, the energy consumption per bit for 100 inferences is $2.73$ times higher than for 1000 inferences. Based on the proposed methodology we also developed an open-source, modular eCAL calculator which enables fine-grained analysis of energy consumption across components, allowing developers to evaluate and optimize AI EE in various deployment scenarios.
Through the proposed eCAL metric, this study lays the foundation for understanding the energy consumption of AI-enhanced emerging communication architectures. Future work may focus on extending this work beyond the assumptions, case studies, and configurations provided in this paper to cover existing and emerging communication system designs that are increasingly relying on AI techniques for their automation and performance optimization, especially the emerging open radio access network (O-RAN) based cellular systems and caching mechanism. The exploration of the relationship between model performance and eCAL for 6G verticals, related network slices, online, distributed, fine tuned and federated model lifecycles may also be worthy of pursuing. Finally, the development of dynamic energy management algorithms in view of a more sustainable operation is highly relevant.
@@ Acknowledgements
This work was supported in part by the HORIZON-MSCA-PF project TimeSmart (No. 101063721), the European Commission NANCY project (No. 101096456), and the Slovenian Research Agency under grants P2-0016, MN-0009-106, and J2-50071.